Operations Report – May 5th

Crew 193 Operations Report 05May2018

SOL over

Name of person filing report: David Attig

Non-nominal systems: Two piece suits still need straps adjusted for neck ring. Generator/inverter/charger. One piece suit #9
Notes on non-nominal systems: The generator still appears to not be charging the batteries at night (only powering the HAB). Shannon is having a look, this should not affect our operations (particularly with the days being so long now).

Generator (hours run): 8:45 pm to 7:30 am.

Solar— SOC 5 % (Before generator is run last night)

Diesel – 60 %

Propane – 60 psi.

ATV Fuel (Ethanol free gasoline) – 0 gallons.

Water (trailer) – 500 gallons.

Water (static) – 500 gallons

Trailer to Static Pump used – no

Water (loft) – Static to Loft Pump used – yes

Water Meter: 134635.2

Toilet tank emptied: no
Date last emptied: May 2

ATV’s (Honda, 300, 350.1, 350.2, 350.3): No

ATV Oil Checked?: No

ATV Oil Added?: No

ATV Fuel Used: 0 Gallons

# Hours the ATVs were Used today: 0

Notes on ATVs: 350.2 is out of service, due to a previous issue. Honda has a hole in the seat. 300 would not start, there is a grinding or clicking sound (dead battery?)

Deimos rover used: No
Hours: 125.6
Beginning charge: 100%
Ending charge: N/A
Currently charging: No

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: Unknown
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Spirit rover used: Yes
Hours: 36.0
Beginning charge: 100 %
Ending charge: 99 %
Currently charging: Yes

Opportunity rover used: Yes
Hours: 31.6
Beginning charge: 100 %
Ending charge: 98 %
Currently charging: Yes

Curiosity rover used: Yes
Hours: 29.9
Beginning charge: 100 %
Ending charge: 98 %
Currently charging: No

Hab Car used?: Yes
Where was Hab Car used?: gas station
Why was Hab Car used?: Water replenishment

General notes and comments: none

Summary of internet: We’re out, same as normal.
Summary of suits and radios: Two piece suits still need strap maintenance

Summary of Hab operations: Super nominal.

Summary of GreenHab operations: Water saturated

Summary of ScienceDome operations: Concluded

Summary of RAM operations: cleaned

Summary of health and safety issues: None to report

Questions, concerns and requests to Mission Support: Sol over

Sol Summary – May 5th

Crew 193 Day 13 Summary Report – 5May18

Day 13

Summary Title: It’s Been A Long Day Without You My Friend, And I’ll Tell You All About It When I See You Again (Alternate Title: Sol Long! Farewell!)

Author: Dr. Shawna Pandya, Crew Health and Safety Officer

Mission Status: Om-nom-nom-nominal

Sol Activity Summary: The day was full and busy as the Commander and XO, no longer bound by duty to travel in separate vehicles, headed toward Earth destination Grand Junction to acquire return vehicles for re-entry to civilization. Meanwhile, the intrepid crew engineer tirelessly plowed on throughout the day on a resupply mission to restock the Hab’s water’s supply. Back at the Hab, the Crew Geologist/Astronomer and Health and Safety officer (wo)manned the fort to buckle down on cleaning duties. The team reunited in the afternoon for a final scrub down of the Hab and surrounding campus, and undertook a final EVA of sorts using an ‘Earth-rover’ (also known as a car) to celebrate with a team dinner. The day can best be summarized as saying, ‘Sol’s well that end’s well…’

Look Ahead Plan:
Back to civilization in the AM!

Anomalies in work: Some unwanted visitors appeared on site, the local authorities were notified.

Weather: Partly cloudy in AM, clear and sunny in the PM.

Crew Physical Status: Nominal, if not a little crunchy on the ol’ bones here in 1g.

A crew EVA was taken up to Marble Ritual on Rovers, for one last team bonding experience. Much merriment ensued.

Reports filed:
Sol Summary

Reports to be filed:

Operations Report

Support Requested: A gentle but loving send-off into the welcoming embrace of Earth’s dawn.

Copyright © The Mars Society. All rights reserved. | Main Site